OpenStudy (anonymous):

solve 0=x3-x2-20x-154.

OpenStudy (loser66):

factor (x-7)(x^2+6x+22) and solve it to get 1 real root and 2 complex ones

OpenStudy (anonymous):

so is the real root 7 then use quadratic to get the other 2

OpenStudy (loser66):

yup

OpenStudy (anonymous):

what did you get when you used quadratic

OpenStudy (loser66):

I don't know, you solve and get it, I can check. It's your stuff, not mine, right? I guide you only. I don't solve it for you

OpenStudy (anonymous):

i know i got (-6+/-√-52)/2

OpenStudy (anonymous):

oh wait it would turn into (-3+/-i√52)/2

OpenStudy (loser66):

yup, simplify them, please . the first one is right, the second one is not

OpenStudy (anonymous):

what do you mean

OpenStudy (loser66):

\[x = \frac {-6\pm\sqrt{-52}}{2}\] is the right answer, but should be simplify and use complex symbol "i" there

OpenStudy (anonymous):

\[x=-3\pm2i \sqrt{13}\]

OpenStudy (loser66):

yup

OpenStudy (anonymous):

thank you so much

OpenStudy (loser66):

nope, sorry, don't have 2 before the sqrt

OpenStudy (loser66):

just \[-3\pm i\sqrt{13}\]

OpenStudy (anonymous):

oh okay